

********************************************************************
UARS WEEKLY STATUS REPORT - 07 AUGUST 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE operations continue normally. We have collected only solar data
this last week during the full sun period. Stellar calibrations experiments
will begin again on Sunday (98221).
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and occulation
data without apparent problems. A yearly calibration is set to begin
this weekend.
***********
PEM
***********
The PEM Zenith boom was deactivated Saturday, August 1. PEM spectrometer
and magnetometer produce atmospheric energy deposition profiles from in situ
electron and proton measurements while they were active.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The MLS team has continued to operate the instrument in a mode for diagnosing
the reoccurance of slips in the antenna scan mechanism, and to attempt to find
a method of operating the instrument which minimizes the impact of these slips
on scientific data.
To date we have not found an operational method which appears acceptable for
routine operation of MLS, and at one point the scan became "stuck" but has now
been freed. We are slowly and carefully pursuing further investigations to
determine how MLS might be operated in an acceptable mode. However, at least
for the current time, MLS cannot be considered operational for obtaining
atmospheric measurements.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 404 CD-ROM requests as of 07-August-1998.
Two orders were received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
105 days HAL (4.80) PROD (L1) (1873R-920C-17F=936L)
28 days HAL (4.80) PROD (L3) (1873R-587C=1286L)
107 days SOL (4.62) PROD (L2-L3) (2349R-942C=1407L)
1 day SEN (5.0) PROD (4R-4C=0L)
2 days SUS (7.0) ENG (2R-2C=0L)
0 days WIN (5.10) ENG (1R-1F=1L)
The CDHF will load June 1998 current day DAAC data in August 1998.
Due to modifications on 07/08/98 to UPAIR node accessibility from DECNet
to IP, a new procedure has been developed by the CDHF and is currently in use
for correlative data ingests from UPAIR.
The CDHF is in the process of reclaiming disk space from special projects
which no longer require these resources. Disk space is currently needed for
Y2K and other development efforts. Current users and UCSS files will not be
effected by this process.
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.5.
***********
FOT
***********
REPORTING PERIOD
This report covers 25 July 1998 (Orbit 37532, GMT Day 206) through 31 July
1998 (Orbit 37635, GMT Day 212).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 210/2225 37605 Off
ACRIM 210/2359 37306 On
SUSIM 210/2359 37306 On
ZEPS 211/0654 37608 On
A summary of Instrument operations supplied by the Mission Planning Group
for the month of July 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle ranged between the angles of 34.6 degrees to 62.4 degrees
for this report period. The beta angle is now increasing toward a maximum
angle of 73.9 degrees on 05 August 1998 (DOY 217).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.8 to 1.3 degrees C. Additional battery performance data
for each battery on 24 July 1998 and 31 July 1998 are provided in Appendix
D.
The clock error ranged from 11.1 to -9.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 05
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% ACRIM 02
% HALOE 01
% MLS 24
% PEM 01
% SUSIM 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/211 show the observatory data loss to be
20 hours, 45 minutes, 42.484 seconds (an increase of 8.736 seconds since
last report period). This is a 0.0346 percent data loss which equals a
99.9654 percent data capture for the mission.
The increase of 8.736 seconds occurred on Day 208 and was due to
insufficient overlap between realtime and recorded data.
Several CMS operational files (one day of FDF data) were accidentally
deleted this week. The deleted files were restored to the CMS from the
daily incremental back-up tape with little impact to the operational
processing timeline.
Applications Processor (AP) ground system software testing has been
suspended waiting release of an operational pack containing version 10.11.
Official delivery of the software was completed on 22 July 1998. The
operational pack will now be generated.
OTHER SIGNIFICANT EVENTS
The B/D thruster refill effort continues. Planning continues and the
engineering thruster tests goal is still 28 July 1998.
The FOT hosted two hands-on CMS training sessions for the benefit of CMS
Year 2K Test Team Lead, Sidney Kirschner (CSC) this week. No actual Year
2K CMS testing has taken place due to the need for FDF products to support
testing. The FOT and Test Team Lead have both requested test planning aids
from FDF, but no products have been delivered.
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the UARS ground system phase-1 re-host effort to
the Lockheed Martin SCS-21 system. Work continued on the test script.
MEETINGS
FOT met NCC representatives on 31 August 1998 to discuss baseline testing
for the NCC portion of the SCS-21 phase 1 demonstration. UARS will use the
NCC-98 system (under test) rather than the operational system for a
demonstration on 05 August 1998. UARS FOT also agreed to validate
performance with the NCC-98 system using the current Application Processor
(AP) at a future date.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 25 August
1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 31 July 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37559 207/2103 98-260 OBC Clk Adj fr 0.00 to 20.6 msec/day
37576 208/2347 98-261 MLS Re-cycle system
37589 209/2031 98-262 MLS Clear Exception Flag
37590 209/2214 98-263 MLS Clear Exception Flag
37591 209/2351 98-264 MLS Clear Exception Flag
37592 210/0137 98-265 MLS MLS Reset
37593 210/0327 98-265 MLS MLS Reset
37595 210/0600 98-265 MLS MLS Reset
37597 210/0918 98-265 MLS MLS Reset
37598 210/1057 98-265 MLS MLS Reset
37599 210/1233 98-265 MLS MLS Reset
37600 210/1414 98-265 MLS MLS Reset
37601 210/1543 98-265 MLS MLS Reset
37602 210/1714 98-265 MLS MLS Reset
37603 210/1905 98-266 OBC Clk Adj fr 20.6 to 0.00 msec/day
37606 211/0047 98-269 ACRIM Power ON
37608 211/0654 98-271 ZEPS Power ON
37610 211/0654 98-268 MLS MLS Reset
37612 211/0924 98-268 MLS MLS Reset
37612 211/0924 98-272 SSPP Op Htrs 1&2 ON BOTH
37613 211/1059 98-268 MLS MLS Reset
37614 211/1237 98-270 SUSIM Power ON
37615 211/1418 98-268 MLS MLS Reset
37616 211/1543 98-268 MLS MLS Reset
37617 211/1715 98-268 MLS MLS Reset
37604 210/2042 98-265 MLS MLS Reset
37605 210/2225 98-267 HALOE Power OFF
37605 210/2225 98-265 MLS MLS Reset
37306 210/2359 98-265 MLS MLS Reset
37306 210/2359 98-269 ACRIM Power ON
37306 210/2359 98-270 SUSIM Power ON
37617 211/1754 98-268 MLS MLS Reset
37617 211/1754 98-273 SUSIM Reloaded RTS 12-Heater Cycle
37618 211/1856 98-274 MLS Microloads
37620 211/2216 98-275 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 31 July 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
QTSISAIF Y-LO SSPP Baseplate by SISA I/F 37532-46
QTSSPPPX Y-HI SSPP +X Thermostat 37540,65,66,72,80,
37581,86,95,602
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37547-617
QTSSPPMX Y-HI SSPP -X Thermostat 37550,72
HRIFESURVTMP Y-HI HRDI IFE Surv HTR TMP 37584
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 31 July 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37559 207/21:03 (26 Jul 98) from 00.0 to 20.6 msec/day
37603 210/19:05 (29 Jul 98) from 20.6 to 00.0 msec/day
37620 211/22:16 (30 Jul 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 31 July 1998
24 June 1998 - GMT Day 205 Beta = 34.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +16.8/ -33.6 +6.76/+4.51 +29.6/-16.0 24.3 1.018 23.0
3 +33.6/ -33.6 +5.47/+3.55 +28.8/-16.0 24.2 1.021 22.4
31 July 1998 - GMT Day 199 Beta = 62.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.1 0.000 0.0
2 +16.8/-22.4 +4.51/+2.92 +22.4/-16.0 26.2 1.025 19.8
3 +44.8/-11.2 +3.23/+1.97 +22.0/-15.2 26.2 1.029 19.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 31 July 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-044 FOT See list July 1998 14170 Generic Late Acquisition
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 31 July 1998
INSTRUMENT SUMMARY FOR MONTH OF July 1998 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- ------------
ACRIM 550 194 0 0
HALOE 524 220 0 0
HRDI 224 520 0 0
MLS 190 554 0 0
PEM (axis) 0 744 0 0
PEM (zeps) 44 700 0 0
PEM (neps) 0 744 0 0
SOLSTICE 744 0 0 0 Note 1
SUSIM 563 181 0 0
WINDII 0 744 0 0
Note 1 - No star data July 24-29
Cold side of spacecraft was facing northward: 14.5 days
Cold side of spacecraft was facing southward: 16.5 days